# Comparing human and LLM politeness strategies in free production

### v1

@@ Acknowledgments
We thank Takateru Yamakoshi for assistance with the LLM evaluation code, Yuka Machino for feedback on the draft, and Noah Goodman for initially thinking this idea would be an interesting one to work on. We also would like to thank anonymous reviewers for providing valuable feedback.
@@ A.2 Additional results
See Table 3 and 4 for comprehensive comparison results between human and LLM responses across different goals. Since the “default” goal case was not studied in Yoon et al. (2020), we focused on “both”, “social”, and “informative” goals and report both Pearson and Spearman correlation scores. For the two tables, we observed that different LLMs have medium to strong correlations with human responses. Both base and instruct-tuned versions of Llama-3.1-8B and Mixtral-8x7B showed very low correlation scores and their incompetence in generating polite language.
See Table 5 and 6 for a complete comparison report between “default” and other goals as a complement to Table 1 in the main text. We reported both Pearson and Spearman correlation scores for “multi-choice-original” and “multi-choice-persona” prompting strategies. We found that including personas does not have any real effect, and the findings are consistent with those reported in the main text.
@@ B.1 Participants
We recruited 156 participants through Prolific in the US or UK to take part in our open-ended response generation task. Participants were compensated at a rate of \$15 / hour following the approved IRB protocol at <University Anonymized>.
@@ B.2 Stimuli
We first needed to elicit a large set of open-ended human responses to compare against the kinds of responses generated by LLMs. To do this, we recruited $N=156$ participants through Prolific, located in the US or UK (compensated at a rate of \$15/hour) and gave them an open textbox to imagine what someone would say in the given scenario. Each participant was assigned 4 distinct scenarios out of the total set of 208 (see Figure 4 middle panel). We planned our sample size to collect at least 3 different responses for each scenario.
@@ C.1 JSD tables
See the complete JSD scores of word-frequency distribution and politeness-strategy distribution at Table 8 and 9.
@@ IV. Bald-on-Record Strategies
C.5 A comprehensive list of politeness strategies with examples
@@
Category Politeness Strategy Example
@@ 18. Be conventionally indirect If you would like …
19. Counterfactual modal forms Could/Would you ……
20. Give deference In my expert opinion, your painting is fabulous.
@@ 34. Use metaphors Your singing was like music to my ears!
Bald-on- Record 35. Direct commands Try practicing with precise measurements.
36. Factuality I didn’t like the cookie at all.
37. Negative lexicon It was terrible

### v2

@@ Acknowledgments
We thank Takateru Yamakoshi for assistance with the LLM evaluation code, Yuka Machino for feedback on the draft, and Noah Goodman for early input and encouragement. We also would like to thank anonymous reviewers for providing valuable feedback.
@@ A.1 Prompting strategies
Multi-choice-original/persona system prompt
@@
NB:
We chose these demographic variables not because we hypothesized that these groups systematically differ in politeness strategies, but rather our use of personas was solely a technique to increase overall response diversity, attempting to approximate the natural variation we see in the population of human responses (as opposed to the modal “original” response). In our analysis, we explicitly did NOT analyze responses by demographic category, as our goal was simply to avoid the kind of mode collapse we observed in preliminary tests without personas.
@@ A.2 Additional results
See Table 4 and 5 for comprehensive comparison results between human and LLM responses across different goals. Since the “default” goal case was not studied in Yoon et al. (2020), we focused on “both”, “social”, and “informative” goals and report both Pearson and Spearman correlation scores. For the two tables, we observed that different LLMs have medium to strong correlations with human responses. Both base and instruct-tuned versions of Llama-3.1-8B and Mixtral-8x7B showed very low correlation scores and their incompetence in generating polite language.
See Table 6 and 7 for a complete comparison report between “default” and other goals as a complement to Table 1 in the main text. We reported both Pearson and Spearman correlation scores for “multi-choice-original” and “multi-choice-persona” prompting strategies. We found that including personas does not have any real effect, and the findings are consistent with those reported in the main text.
@@ B.1 Participants
We recruited 156 participants through Prolific in the US or UK to take part in our open-ended response generation task. Participants were compensated at a rate of \$15 / hour following the IRB protocol.
B.1.1 Participant Demographics
Stage US UK Female Male
Open-ended Response
Collection (n=156)
80.1% 19.9% 59.7% 40.4%
Human Evaluation
(n=156)
75.1% 24.9% 58.8% 40.7%
Table 9: Participant Demographics
See Table 9 for specific demographics of participants in our human studies. To capture the diversity of our participant pool, we also collected self-reported ethnicity information. During the open-ended response collection stage, most participants identified as White (63.28%), with the remainder identifying as Black (18.64%), Mixed (7.91%), Asian (5.08%), or Other ethnic backgrounds (5.08%). In the human evaluation stage, the distribution shifted slightly, with 57.31% identifying as White, 18.71% as Black, 10.53% as Asian, 5.85% as Mixed, and 7.02% as Other ethnic backgrounds.
@@ B.2 Stimuli
We first needed to elicit a large set of open-ended human responses to compare against the kinds of responses generated by LLMs. To do this, we recruited $N=156$ participants through Prolific, located in the US or UK (compensated at a rate of \$15/hour) and gave them an open textbox to imagine what someone would say in the given scenario. Each participant was assigned 4 distinct scenarios out of the total set of 208 (see fig:experiment-pipeline middle panel). We planned our sample size to collect at least 3 different responses for each scenario.
@@ B.5 Prompting strategy
B.6 Calibration on LLMs’ open-ended generation
To calibrate the choice of models in Experiment 2, we ran the open-ended experiment with Llama-3.1-8B-Instruct, one of the smaller models from Experiment 1, using an identical experimental setup. We collected 208 responses, one for each scenario. The results revealed severe limitations in the smaller model’s capabilities. Despite receiving identical prompts (instructing concise responses), Llama-3.1-8B exhibited multiple failure modes: (1) it generated excessively long responses or incomplete sentences, (2) failed to follow basic instructions, and (3) frequently hallucinated content unrelated to the scenarios.
For example, when asked to provide informative feedback for a 1-heart rating, the model generated the following hallucinated ratings explanations:
Don’t try to be too different. Your presentation was good, but could have been smoother." Rating: 0 hearts (very direct).
This would not be true because Bob just wanted general feedback as opposed to negative feedback. Rating: 2 hearts (relatively kind but still true) (0, this might hurt Bob’s feelings; 3, perfect)
In more extreme cases, responses devolved into random character sequences after one or two coherent sentences:
That was one of the least, uh…clear presentations I’ve heard in a while. However, I really like your confidence on stage, it’s something you can work with! “ F A S L F B H M B F I S S M T B U S P K E Y T O I E N S P A S S I N G D E L I G H T E S S I G.”
To enable a meaningful comparison with models used in the main analysis, we truncated all Llama-3.1-8B responses to the first complete sentence and annotated them for politeness strategies using the same methodology as our main analysis. The resulting distributions differed significantly from all larger models tested, with Jensen-Shannon Divergence values exceeding 0.53 ($p<0.001$) in all comparisons. These findings suggest that model size represents a critical threshold for pragmatic competence, even if we ignore basic failures to follow task instructions. While we cannot isolate whether this stems from training data, reinforcement strategies, or architectural limitations, this calibration experiment confirms our initial decision to focus on models that demonstrate basic pragmatic competence. Additionally, the investigation of size thresholds will be an important direction of future work.
@@ C.1 JSD tables
See the complete JSD scores of word-frequency distribution and politeness-strategy distribution at Table 10 and 11.
@@ IV. Bald-on-Record Strategies
C.5 Guidelines for correcting LLM annotation results
We calculated the initial agreement between our two LLM annotators (GPT-4.1 and Claude-3.7-Sonnet) at 39.8% (using a strict measure requiring an exact match of the set of strategies), indicating substantial divergence in their initial classifications. When compared to the human-corrected “gold standard” labels, Claude-3.7-Sonnet achieved 34.3% agreement while GPT-4.1 achieved 36.1% agreement. Because each utterance could receive multiple strategy labels, we observed that many LLM annotations contained a mix of appropriate and inappropriate labels. Rather than rejecting these labels entirely, we systematically corrected them by removing inappropriate labels and retaining valid ones. The authors first established the correction criteria through discussion, then independently reviewed annotations.
We identified four main categories of LLM annotation errors:
  1. 1.
    LLM annotators often labelled strategies without considering the scenario context that we provided. For instance, “there is still room for improvement” was labeled as an “understatement” regardless of the true rating. This label would be appropriate for poor ratings (0/3 hearts) but inappropriate when the true rating was already positive (2/3 hearts).
  2. 2.
    Relatedly, when ratings literally matched the content (e.g., “really good” for a 3-heart rating), LLMs annotators sometimes incorrectly tagged them as “exaggeration” strategies.
  3. 3.
    In about 20 cases, LLMs generated strategies not present in the prompt, such as “Negative Politeness – Be specific”, which doesn’t exist in the taxonomy of Brown and Levinson (1987).
  4. 4.
    In a remaining 5-10% of cases, annotations were entirely wrong, such as labeling “that really makes a solo shine” as a hedge when it clearly expresses a positive evaluation.
C.6 A comprehensive list of politeness strategies with examples
@@
Category Specific Politeness Strategy Example
@@ 18. Be conventionally indirect If you would like …
19. Counterfactual modal forms Could/Would you …
20. Give deference In my expert opinion, your painting is terrible.
@@ 34. Use metaphors Your singing was like music to my ears!
Bald on- Record 35. Direct commands Try practicing with precise measurements.
36. Factuality I didn’t like the cookies at all.
37. Negative lexicon It was terrible!